the population of a small town will double every 5 years the population is 47 in 2005 what will the population be in 2030
1
Expert's answer
2012-10-12T08:41:36-0400
During 2005-2030 the population will double

N = (2030 - 2005)/5 = 5

times. So, in 2030 the population will be

M = 47*2^5 = 1504.